# Modern Hopfield Networks meet Encoded Neural Representations - Addressing Practical Considerations

### v1

@@ 2.1 Equivalence with Kernel Memory Networks (KMN)
The formulation of the MHNs can be viewed as a special case of Kernel Memory Networks (KMNs) (Iatropoulos et al., 2022), where each neuron performs kernel-based classification or regression. The MHN update rule is analogous to that in KMNs and can be computed in closed form as a recurrence.
Let ${K}(\mathbf{\Xi},\mathbf{s}^{(t)})$ denote the (symmetric positive definite) kernel function that defines pairwise similarities, and $\mathbf{K}^{\dagger}$ be the Moore-Penrose pseudoinverse for $\mathbf{K}(\mathbf{\Xi},\mathbf{\Xi})$. For continuous valued memories, the radial translation-invariant exponential kernel (infinite dimensional basis) with a fixed spatial scale $r$ and temperature $\alpha$ is proposed. This parameterization allows for an analysis of memory capacity and storage limits by interpreting the MHN optimization as a feature transformation operating in a Reproducing Kernel Hilbert Space (RKHS). The kernel and state update are as follows:
@@
While KMNs provide strong theoretical storage guarantees (Iatropoulos et al., 2022), their real-world performance can be sensitive to data distributions and parameter choices.
@@ 2.2 Equivalence with Transformers
The transformer kernel can be written as: $K(\mathbf{x},\mathbf{y})=\exp{\Big{[}\frac{1}{\sqrt{d_{k}}}(\mathbf{W}_{Q}\mathbf{x})^{T}(\mathbf{W}_{K}\mathbf{y})\Big{]}}$. Overall, this highlights the connection between the MHNs and transformer attention mechanisms, showing that both rely on projecting input representations into a shared space for similarity-based comparison.Further details on the KMN and transformer equivalence are provided in Appendix Section A.
@@ 3.1 Experimental Evaluation and Results:
Specifically, we evaluated various pre-trained encoder-decoder architectures known for their state-of-the-art performance in deep learning-based image encoding and decoding ranging from vanilla-transformer based to variational auto-encoder based models. In particular, we utilized the Discrete Variational Autoencoder (D-VAE) from (Ramesh et al., 2021) and other architectures from (Rombach et al., 2021) and explored two transformer variants from the diffusion library: one trained with codebook-based (Vector Quantized - VQ) criteria and the other using Kullback-Leibler (KL) divergence-based criteria. Our empirical analysis revealed that Vector Quantized VAE (VQ-VAE) methods outperformed others in our setup. Consequently, we selected D-VAE and variants of VQ-F8 and VQ-F16 from (Rombach et al., 2021) for further analysis. To maintain consistency in representation, we downsampled all the images to a resolution of $28\times 28\times 3$ to match the number of features that the encoded representations produced.
@@
KMNs Sensitivity to Real-World Conditions: KMNs have strong theoretical guarantees, but in our experiments, they faced significant challenges with real-world, large-scale image retrieval tasks. Despite an extensive parameter sweep over hyperparameters like $r$ (spatial scale) and $\alpha$ (inverse temperature), KMNs were highly sensitive to these settings and consistently underperformed. For example, with a memory bank of 6000 images, KMNs yielded an MSE of 0.2427 and 1-SSIM of 0.972, indicating poor retrieval accuracy. While KMNs excel in controlled environments, their performance degrades in high-dimensional, non-linearly separable data, limiting their practical use. In contrast, our HEN shows more robust performance, mitigating meta-stable states and improving retrieval accuracy.
@@
Scale-out performance: We studied the performance of the HEN on an increasing number of images in the memory bank. Table 1 shows the performance metrics across different encoder approaches with an increasing number of images. All encoder-based approaches robustly recover the encoder image representations without seeing a drop in the quality of the reconstruction with an increasing number of images stored in $\{\mathbf{\xi}_{n}\}$. Section 2 of the Appendix contain a more detailed breakdown of the recovery performance as a function of encoder-decoder methods, and the similarity metric for different values of $\beta$. By contrast, KMNs consistently failed to scale with the increase in images, showing degraded performance as the number of images increased.
@@
Limitations: While our approach demonstrates robust performance, we do not claim a universal guarantee that any combination of time step $T_{f}$, update rule, or encoder-decoder methods will always yield perfect convergence. Our results indicate that after a certain number of iterations ($T_{f}=100$ in our case), the dynamics reach a plateau, and the optimal $T_{f}$ can vary depending on the dataset. Thus, while our method delays the onset of metastable states, performance may vary in other contexts.
@@ 3.2 Natural language-based Hetero-associations
Figure 6: In each example, (Left) the upper and lower image cases depict the disruption of unique associations in the $\{\mathbf{\xi}_{n}\}$ memory bank. (Right) A single text input corresponding to these disrupted associations is used during the query phase. The top blue image represents an empty image as a zero-encoded vector at initialization. The reconstruction appears to be a meta-stable state. Neither of the original images is accurately recovered, supporting the hypothesis of unique text-image associations.
@@ 4 Conclusions
In this paper, we began by unifying and exploring various energy formulations used across associative memory models, highlighting their interconnections and the common emphasis on pattern separability. This theoretical foundation informed our work on improving Modern Hopfield Networks.
Building on these insights, we introduced two key enhancements: pattern encoders and decoders integrated with Modern Hopfield Networks to enhance pattern separability and reduce metastable states. Additionally, we demonstrated how this approach supports cross-stimuli associations using different encodings, as long as the uniqueness of association is maintained. These advancements mark a step toward improving the practicality and performance of Modern Hopfield Networks for real-world retrieval and storage tasks.
@@
satyananda.kashyap@ibm.com Niharika S. D’Souza
IBM Research - Almaden Luyao Shi
IBM Research - Almaden Ken C. L. Wong
IBM Research - Almaden Hongzhi Wang
IBM Research - Almaden Tanveer Syeda-Mahmood
@@ A.1 Relating Transformer and HEN Updates to the KMN formulation:
Under the KMN formulation, the MHN from involves a kernel that is not symmetric positive definite, as the energy objective is not a convex optimization problem (Iatropoulos et al., 2022; Wright & Gonzalez, 2021). Nevertheless, it permits a bilinear reproducing form for the kernel $K(\cdot,\cdot)$, where the feature transformation maps $\{\mathbf{\Phi}_{\mathcal{X}}(\cdot):\mathcal{X}\rightarrow\mathcal{F}_{\mathcal{X}}(\cdot),\mathbf{\Phi}_{\mathcal{Y}}(\cdot):\mathcal{Y}\rightarrow\mathcal{F}_{\mathcal{Y}}\}$ map input elements from vector spaces $\{\mathbf{x}\in\mathcal{X},\mathbf{y}\in\mathcal{Y}\}$ to the D-dimensional Banach spaces $\{{\mathbf{\Phi}_{\mathcal{X}}}(\cdot)\in\mathcal{F}_{\mathcal{X}},{\mathbf{\Phi}_{\mathcal{Y}}}(\cdot)\in\mathcal{F}_{\mathcal{Y}}\}$, characterized by linear manifold functional forms $\{\mathbf{f}_{\mathbf{\Phi}_{\mathcal{X}}}(\mathbf{x};\mathbf{W}_{Q})\in\mathcal{B}_{\mathcal{X}},\mathbf{g}_{\mathbf{\Phi}_{\mathcal{Y}}}(\mathbf{y};\mathbf{W}_{K})\in\mathcal{B}_{\mathcal{Y}}\}$. The mathematical expression for the transformer kernel is:
@@
where $C$ is a constant. This is the form of the transformer kernel corresponding to the update in Eq. (2) and energy functional in Eq. (1) of the main manuscript. We note that spherical normalization was also used in  (Ramsauer et al., 2020; Krotov & Hopfield, 2020) as a mechanism to control the dynamics of the updates to mitigate metastable solutions.
@@
Additionally, delaying the onset of metastable states is an important feature of our proposed approach, as it enables the system to store more content within encoded representations, potentially leading to a higher memory-capacity ceiling. However, this does not imply that the method is entirely free from metastable states or that it will always guarantee convergence for all possible input distributions. Certain real-world datasets with high noise, irregular structure, or non-separable patterns may still present challenges, even for our enhanced architecture.
Future work could involve extending the analysis of metastable state dynamics across different datasets and exploring further variations of update rules or encoding strategies to improve convergence guarantees across a wider range of practical applications.
@@ Appendix D Handling cross-stimuli using pixelized language-image association
To test language-image associations, we utilized the unique set of captions associated with each image in the COCO dataset. Specifically, we employed Python’s hashlib.sha256(.) function to hash the captions generating a unique ID text string to associate with the image. Initially, we created a memory bank $\{{\mathbf{\xi}}^{\bf T}_{n}\}$ by converting the hashed captions into pixelized text representations using a generic text-to-pixel function. Subsequently, both the pixelized text and the corresponding images were processed through the same encoder. The resulting vectors were concatenated to form the elements of the memory bank $\{\hat{\mathbf{\xi}}_{n}=[\hat{\mathbf{\xi}}^{\bf I}_{n};{\mathbf{\xi}}^{\bf T}_{n}]\}$.

### v2

@@ 2.1 Equivalence with Kernel Memory Networks (KMN)
The formulation of the MHNs can be viewed as a special case of Kernel Memory Networks (KMNs) (Iatropoulos et al., 2022), where each neuron performs kernel-based classification or regression. The MHN update rule is analogous to that in KMNs and can be computed in closed form as a recurrence. Let ${K}(\mathbf{\Xi},\mathbf{s}^{(t)})$ denote the (symmetric positive definite) kernel function that defines pairwise similarities, and $\mathbf{K}^{\dagger}$ be the Moore-Penrose pseudoinverse for $\mathbf{K}(\mathbf{\Xi},\mathbf{\Xi})$. For continuous valued memories, the radial translation-invariant exponential kernel (infinite dimensional basis) with a fixed spatial scale $r$ and temperature $\alpha$ is proposed. This parameterization allows for an analysis of memory capacity and storage limits by interpreting the MHN optimization as a feature transformation operating in a Reproducing Kernel Hilbert Space (RKHS). The kernel and state update are as follows:
@@
While KMNs provide strong theoretical storage guarantees (Iatropoulos et al., 2022), their real-world performance is known to be sensitive to data distributions and parameter choices Wu et al. (2024).
@@ 2.2 Equivalence with Transformers
The transformer kernel can be written as: $K(\mathbf{x},\mathbf{y})=\exp{\Big{[}\frac{1}{\sqrt{d_{k}}}(\mathbf{W}_{Q}\mathbf{x})^{T}(\mathbf{W}_{K}\mathbf{y})\Big{]}}$. Overall, this highlights the connection between the MHNs and transformer attention mechanisms, showing that both rely on projecting input representations into a shared space for similarity-based comparison. Further details on the KMN and transformer equivalence are provided in Appendix Section 1.
@@ 3.1 Experimental Evaluation and Results:
Specifically, we evaluated various pre-trained encoder-decoder architectures known for their state-of-the-art performance in deep learning-based image encoding and decoding ranging from vanilla-transformer based to variational auto-encoder based models. In particular, we utilized the Discrete Variational Autoencoder (D-VAE) from (Ramesh et al., 2021) and other architectures from (Rombach et al., 2021) and explored two transformer variants from the diffusion library: one trained with codebook-based (Vector Quantized - VQ) criteria and the other using Kullback-Leibler (KL) divergence-based criteria. Our empirical analysis revealed that Vector Quantized VAE (VQ-VAE) methods outperformed others in our setup. Consequently, we selected D-VAE and variants of VQ from (Rombach et al., 2021) for further analysis. To maintain consistency in representation, we downsampled all the images to a resolution of $28\times 28\times 3$ to match the number of features that the encoded representations produced.
@@
KMNs Sensitivity to Real-World Conditions: KMNs have strong theoretical guarantees, but in our experiments, they faced significant challenges with real-world, large-scale image retrieval tasks. Despite an extensive parameter sweep over hyperparameters like $r$ (spatial scale) and $\alpha$ (inverse temperature), KMNs were highly sensitive to these settings and consistently underperformed. For example, with a memory bank of 6000 images, KMNs yielded an MSE of 0.2427 and 1-SSIM of 0.972, indicating poor retrieval accuracy. While KMNs excel in controlled environments, their performance degrades in high-dimensional, non-linearly separable data, limiting their practical use. This is suspected to be a direct consequence of the restrictive assumptions made on the underlying data-distributions in KMNs (Wu et al., 2024). In contrast, our HEN shows more robust performance, mitigating meta-stable states and improving retrieval accuracy.
@@
Scale-out Performance and Ablation Study on Encoder-Decoder Pairs: We evaluated HEN’s scalability by testing it’s retrieval performance as the number of stored images increased in the memory bank. Table 1 presents the performance metrics across different encoder-decoder approaches as the memory bank scaled from 6,000 to 15,000 images. Our results show that all the encoder-based approaches robustly recovered the image representations without a noticeable drop in the reconstruction quality, supporting HEN’s stability and scalability in retrieval performance at this scale. Additionally, our ablation studies included five pre-trained encoder-decoder architectures - dVAE  (Ramesh et al., 2021) and VQ-F8, VQ-F16, KL-F8, KL-F16 from (Rombach et al., 2021) to assess whether HEN’s retrieval stability holds across varied configurations. Section 2 of the Appendix contain a more detailed breakdown of the recovery performance as a function of encoder-decoder methods, and the similarity metric for different values of $\beta$. By contrast, KMNs consistently failed to scale with the increase in images, showing degraded performance as the number of images increased.
@@ 3.2 Natural language-based Hetero-associations
Figure 6: In each example, (Left) the upper and lower image cases depict the disruption of unique associations in the $\{\hat{\mathbf{\xi}}_{n}\}$ memory bank. (Right) A single text input corresponding to these disrupted associations is used during the query phase. The top blue image represents an empty image as a zero-encoded vector at initialization. The reconstruction appears to be a meta-stable state. Neither of the original images is accurately recovered, supporting the hypothesis of unique text-image associations.
@@ 4 Conclusions
In this paper, we unified and explored the diverse energy formulations used across associative memory models, highlighting their theoretical interconnections and the shared focus on pattern separability. This foundational analysis informed our development of key enhancements for Modern Hopfield Networks (MHNs), specifically by integrating pattern encoders and decoders to enhance separability and reduce metastable states. Additionally, we demonstrated how this approach supports cross-stimuli associations using different encodings, as long as the uniqueness of association is maintained suggesting promising adaptability for applications that require cross-modal recall, like multimedia search or multi-sensor data fusion. These advancements mark a step toward improving the practicality and performance of Modern Hopfield Networks for real-world retrieval and storage tasks.
Additionally, HEN’s architecture shows potential for other associative tasks, such as temporal sequences or visual variations (e.g., images of the same entity from different angles). Prior studies on associative memory Gutfreund & Mezard (1988); Chaudhry et al. (2023); Shriwas et al. (2019); Millidge et al. (2022) suggest that with minor modifications, HEN could be extended to sequential or contextual retrieval, supporting broader applications in dynamic and context-rich environments.
While HEN has shown consistent performance, we recognize that its convergence and stability depend on factors such as the update rule, encoder-decoder selection, and time step $T_{f}$. Our results indicate that after a set number of iterations ($T_{f}=100$ in our case), retrieval dynamics reach a stable plateau; however, the optimal $T_{f}$ may vary by dataset. Thus, while our approach delays metastable states, further tuning may be necessary in different contexts.
Looking forward, future work could investigate HEN’s scalability with even larger datasets by implementing fine-tuning steps to optimize VAE parameters, potentially further reducing metastable behavior. This approach could enhance HEN’s memory capacity, enabling it to support practical storage and retrieval use cases at scale. These extensions position HEN as a versatile framework, adaptable to diverse associative memory tasks and scalable to meet the demands of large-scale, heterogeneous data environments.
@@
satyananda.kashyap@ibm.com &Niharika S. D’Souza
IBM Research - Almaden &Luyao Shi
IBM Research - Almaden &Ken C. L. Wong
IBM Research - Almaden &Hongzhi Wang
IBM Research - Almaden &Tanveer Syeda-Mahmood
@@ A.1 Relating Transformer and HEN Updates to the KMN formulation:
Under the KMN formulation, the MHN form involves a kernel that is not symmetric positive definite, as the energy objective is not a convex optimization problem (Iatropoulos et al., 2022; Wright & Gonzalez, 2021). Nevertheless, it permits a bilinear reproducing form for the kernel $K(\cdot,\cdot)$, where the feature transformation maps $\{\mathbf{\Phi}_{\mathcal{X}}(\cdot):\mathcal{X}\rightarrow\mathcal{F}_{\mathcal{X}}(\cdot),\mathbf{\Phi}_{\mathcal{Y}}(\cdot):\mathcal{Y}\rightarrow\mathcal{F}_{\mathcal{Y}}\}$ map input elements from vector spaces $\{\mathbf{x}\in\mathcal{X},\mathbf{y}\in\mathcal{Y}\}$ to the D-dimensional Banach spaces $\{{\mathbf{\Phi}_{\mathcal{X}}}(\cdot)\in\mathcal{F}_{\mathcal{X}},{\mathbf{\Phi}_{\mathcal{Y}}}(\cdot)\in\mathcal{F}_{\mathcal{Y}}\}$, characterized by linear manifold functional forms $\{\mathbf{f}_{\mathbf{\Phi}_{\mathcal{X}}}(\mathbf{x};\mathbf{W}_{Q})\in\mathcal{B}_{\mathcal{X}},\mathbf{g}_{\mathbf{\Phi}_{\mathcal{Y}}}(\mathbf{y};\mathbf{W}_{K})\in\mathcal{B}_{\mathcal{Y}}\}$. The mathematical expression for the transformer kernel is:
@@
where $C$ is a constant. This is the form of the transformer kernel corresponding to the update in Eq. (2) and energy functional in Eq. (1) of the main manuscript. We note that spherical normalization was also used in  (Ramsauer et al., 2020; Krotov & Hopfield, 2020; Wu et al., 2024) as a mechanism to control the dynamics of the updates to mitigate metastable solutions.
@@
Additionally, delaying the onset of metastable states is an important feature of our proposed approach, as it enables the system to store more content within encoded representations, potentially leading to a higher memory-capacity ceiling. However, this does not imply that the method is entirely free from metastable states or that it will always guarantee convergence for all possible input distributions.
Certain real-world datasets with high noise, irregular structure, or non-separable patterns may still present challenges, even for our enhanced architecture. Future work could involve extending the analysis of metastable state dynamics across different datasets and exploring further variations of update rules or encoding strategies to improve convergence guarantees across a wider range of practical applications.
@@ Appendix C Quantifying Meta-Stable States
Finally, the work of (Wu et al., 2024) leverages the equivalence between MHNs, kernel methods, and transformers to directly address the issue of separability of patterns under arbitrary data-distributions. Their framework introduces a second separation maximization objective that is added to the energy minimization objective (i.e. Eq.(1) in the main manuscript). The separation objective is based on a kernel transformation parameterized as a learnable linear kernel or deep network (i.e. encoder) and radial kernel. Thus, pattern retrieval is a two step process that interleaves a stochastic optimization of the encoding parameters with the retrieval dynamics of the Modern Hopfield Network. Here, in each step, local convergence is achieved either as a mini-batch or full batch gradient descent routine. This strategy was shown to provide improvements over the traditional MHNs and their variants upto dataset sizes of 500 for image retrieval. In contrast, our framework fixes the encoding transformation as defined by a pre-trained VAE. As the VAE parameters are frozen apriori, this avoids an expensive outer optimization loop, in turn providing expedient retrieval for datasets of relatively larger size (15000 examples) as well as support hetero-association.
@@ Appendix D Handling cross-stimuli using pixelized language-image association
To test language-image associations, we utilized the unique set of captions associated with each image in the COCO dataset. Specifically, we employed Python’s hashlib.sha256(.) function to hash the captions generating a unique ID text string to associate with the image. Initially, we created a memory bank $\{{\hat{\mathbf{\xi}}}^{\bf T}_{n}\}$ by converting the hashed captions into pixelized text representations using a generic text-to-pixel function. Subsequently, both the pixelized text and the corresponding images were processed through the same encoder. The resulting vectors were concatenated to form the elements of the memory bank $\{\hat{\mathbf{\xi}}_{n}=[\hat{\mathbf{\xi}}^{\bf I}_{n};{\hat{\mathbf{\xi}}}^{\bf T}_{n}]\}$.